he angle between two vectors a and bis 2pi/3 and modulus a=1,modulus b=2 Evaluate Mod a cross3b 

Grade:12

1 Answers

BALAJI ANDALAMALA
8 years ago
|a\times 3b|=3|a||b|sin\theta
=3\times2\times sin120
=3\times2\times\frac{\sqrt{3}}{2} = 3\sqrt{3}